# Taxonomy and Estimation of Multiple Breakpoints in High-Dimensional Factor Models

### v2

@@
Abstract:
This paper proposes a Quasi-Maximum Likelihood (QML) estimator for break points in high-dimensional factor models, specifically accounting for multiple structural breaks. We begin by establishing a necessary and sufficient condition to categorize two distinct types of breaks in factor loadings: singular changes and rotational changes. By examining the properties of nearly singular subsample pseudo factor covariance matrices, we show that the QML estimator precisely identifies the true breakpoint with probability tending to one in large samples for singular changes. For rotational changes, we demonstrate that the estimator achieves estimation errors that are asymptotically bounded by $O_{p}(1)$ (fraction consistency). Furthermore, we introduce an information criterion (IC) to estimate the number of breaks, proving that it can detect the true number with probability tending to one in large samples. Monte Carlo simulations confirm the strong finite sample performance of our proposed methods. Finally, we provide an empirical example to estimate structural breakpoints in the FRED-MD dataset spanning 1959 to 2024.
Key words and phrases: Quasi-Maximum Likelihood, Singularity, Consistency, Information criterion.
@@ 1 Introduction
Factor models have become a cornerstone in the analysis of large-scale economic and financial datasets characterized by both high cross-sectional and time dimensions. The models’ versatility has led to widespread applications across economics and statistics, including the study of policy shocks (Stock and Watson, 2016; Han, 2023), macroeconomic dynamics (Boivin and Giannoni, 2006), asset pricing (Giglio and Xiu, 2021; Giglio et al., 2025), portfolio theory (Fan et al., 2008, 2013; Ding et al., 2021), causal inference (Gobillon and Magnac, 2016; Xu, 2017), and high-dimensional statistical machine learning (Fan et al., 2021). Yet, despite their wide applicability, real-world data are rarely stable over time. Structural breaks are ubiquitous, arising from policy interventions, financial crises, technological progress, or global events such as the COVID-19 pandemic. Importantly, they often embody the causal impact of interventions and shocks-policies may reshape economic structures, innovations may alter production and financial linkages, and major crises may transform underlying relationships. Identifying and modeling these breaks is therefore a powerful tool for causal inference, policy evaluation, and understanding how economic systems evolve in response to major events.
@@
Previous studies (e.g., Han and Inoue, 2015) show that a single structural break in the factor loading matrix admits an observationally equivalent representation with time-invariant loadings and an expanded set of pseudo-factors. This expansion induces singularity in the covariance matrices of the pseudo-factors from either the pre-break or post-break subsample. Duan, Bai, and Han (2023, hereafter DBH) exploit this property and demonstrate that such singularity is key to establishing the consistency of the QML estimator in the single-break setting. However, the DBH approach is restricted by its single-break assumption, even though multiple breakpoints commonly arise in practice.
@@
Partly due to the simultaneous existence of both singular and rotational breaks, extending the analysis from a single-break to a multi-break setting presents several significant and non-intuitive challenges. First, unlike the single-break case, merely checking the singularity of the pseudo-factor covariance matrix within a given regime is insufficient to characterize the type of break in a multi-break environment. A break might appear rotational even if neither the preceding nor the succeeding regime exhibits full column rank, a phenomenon that can occur due to singular breaks at other time points. Thus, a primary challenge lies in rigorously defining rotational and singular changes in factor loadings within a multi-break framework. Second, because multiple breakpoints are jointly estimated, the estimation errors of these individual breakpoint estimators become inherently interdependent, posing substantial technical challenges for asymptotic analysis. Third, given that the QML estimator exhibits distinct convergence rates for singular and rotational breaks, consistently determining the true number of breaks without prior knowledge of their specific types emerges as a challenging task.
@@
We address these challenges and make the following contributions. First, we establish a necessary and sufficient condition to precisely distinguish between singular and rotational structural breaks in factor loadings within a multi-break framework. Specifically, for each breakpoint, break types are identified by comparing the numbers of effective factors in the pre-break and post-break regimes with the number of effective factors in the regime formed by pooling these two regimes. Second, to tackle the interdependence of estimation errors mentioned previously, we analyze the asymptotic properties of the $j$-th breakpoint by overfitting the remaining true breakpoints. This strategy requires us to compare the eigenvalues of the sample covariance matrix of the pseudo-factors in a regime against those from a subsample of that same regime. Third, leveraging this crucial result, we prove that the QML estimator identifies the true breakpoint with probability tending to one in large samples for singular breaks, and achieves $O_{p}(1)$ estimation errors for rotational breaks. Finally, we analyze how underfitting and overfitting the number of breaks affects the value of our QML objective function under both types of breaks. This analysis enables us to design an effective penalty function for our information criterion, ensuring its consistent selection of the true number of breaks without requiring prior knowledge of break types.
@@
Links to the literature. A growing body of literature investigates tests for structural breaks in high-dimensional factor models (e.g., Chen et al., 2014; Han and Inoue, 2015; Yamamoto and Tanaka, 2015; Bai et al., 2024; Fan et al., 2024; Peng et al., 2025). Our study differs by focusing on the estimation of breakpoints rather than testing for their presence. Moreover, traditional literature on structural break estimation primarily addresses time series with small cross-sectional dimensions (e.g., Bai and Perron, 1998; Bai, 2000; Qu and Perron, 2007; Chan et al., 2014, 2021). In contrast, our factor model framework operates in a setting where both the cross-sectional ($N$) and time ($T$) dimensions are large.
@@
Within this large $N$, large $T$ context, prior work has explored structural breaks in various settings. Qian and Su (2016), for instance, applied group fused Lasso to panel models with structural breaks, and Li et al. (2016) extended this to panel data with interactive fixed effects. While these studies consider large panels, their focus differs from ours, as they concentrate on breakpoints in regression coefficients rather than the high-dimensional factor loading matrix. For high-dimensional time series, Bai (2010) developed a Least Squares (LS) estimator for breakpoints in the mean and variance, and Li et al. (2024) introduced a two-way MOSUM procedure with asymptotic properties for change-point estimators in trend functions. Our framework distinguishes itself from these by specifically addressing breakpoints in the loading coefficients of unobserved factors.
@@
Regarding breakpoint estimation in factor models, earlier literature established the consistency of the estimated break fraction, specifically that $\hat{\tau}-\tau_{0}=o_{p}(1)$, where $\hat{\tau}$ is the estimated breakpoint $\hat{k}$ divided by sample size $T$, and $\tau_{0}$ is the true fraction (e.g., Chen, 2015; Cheng et al., 2016). More recent studies have provided theoretical results on the detailed rate of the error bound for the estimated breakpoint. For example, Barigozzi et al. (2018) proposed a wavelet-based method to estimate both the number and locations of structural breaks in common and idiosyncratic components. Barigozzi et al. (2024) further developed a MOSUM procedure for detecting multiple change points in high-dimensional factor models. Both studies showed that the estimation error of the breakpoint diverges slowly with sample size, implying a consistently estimated break fraction. Similarly, Baltagi et al. (2017, 2021) developed an estimator that compares subsample second moments of factors estimated via full-sample principal components. Their breakpoint estimator exhibits a stochastically bounded estimation error, which also implies a consistently estimated break fraction.
@@
In contrast to these approaches, our QML estimator achieves not only fraction-consistency but also point-consistency, meaning the estimated breakpoints equal the true ones with probability approaching one in large samples. To our knowledge, only a few studies have established point-consistency for breakpoint estimators of the loading matrix in factor models. Bai et al. (2020) demonstrated that the LS estimator is point-consistent under both large and small break settings. However, their framework is limited to a single break, whereas our methodology accommodates multiple breaks. Ma and Su (2018) proposed a fused Lasso estimator that is consistent for multiple breakpoints in factor models, under the assumption of a constant number of factors. Our work considers a more general and empirically relevant setting by allowing the number of effective factors to vary over time. Additionally, Ma and Tu (2023) adopted a group fused Lasso approach to estimate changes in regression coefficients of one factor on the remaining factors. Their fused Lasso estimator exhibits stochastically bounded errors, which are then refined by LS to achieve point-consistency. Nonetheless, their method does not accommodate certain types of rotational changes (e.g., when the post-break loading matrix is rescaled by a constant), a feature our Type 2 break in Section 3 does allow.
@@
Terminology and Notation. Throughout the paper, we will use the concept of the number of pseudo factors. While we will define it more precisely after introducing the model, it generally refers to the number of factors needed to capture the common variations in the data. This concept can be applied to the entire sample or to specific subsamples (e.g., a subsample that includes two adjacent regimes). In each case, it denotes the number of factors required to explain the common variation within the sample or subsample. Thus, the number of pseudo factors for the whole sample is no smaller than that of a subsample.
The following notations will be used throughout the paper. Let $\rho_{i}(\mathbb{B})$ denote the $i$-th eigenvalue of an $n\times n$ symmetric matrix $\mathbb{B}$, and $\rho_{1}(\mathbb{B})\geq\rho_{2}(\mathbb{B})\geq\cdots\geq\rho_{n}(\mathbb{B})$. For an $m\times n$ real matrix $\mathbb{A}$, we denote its Frobenius norm as $\|\mathbb{A}\|=[tr(\mathbb{A}\mathbb{A}^{{}^{\prime}})]^{1/2}$, and its adjoint matrix as $\mathbb{A}^{\#}$ when $m=n$. Let $\mathrm{Proj}(\mathbb{A}|\mathbb{Z})$ denote the projection of matrix $\mathbb{A}$ onto the columns of matrix $\mathbb{Z}$. For a,b>0a,\ b>0, $a\gtrsim b$ indicates that there exists a constant $c>0$ such that $a\geq cb$. We write $a\asymp b$ to mean that there exist constants $c_{1},c_{2}>0$ such that $c_{1}b\leq a\leq c_{2}b$. For a real number $x$, $[x]$ represents the integer part of $x$.
@@ 2.1 Model
Let us consider the following factor model with $m_{0}$ common breaks at $k_{1}^{0}(T)<k_{2}^{0}(T)<\cdots<k_{m_{0}}^{0}(T)$ in the factor loadings for $i=1,\cdots,N$:
$x_{it}=\begin{cases}\lambda_{i,1}f_{t}+e_{it}&for\leavevmode\nobreak\ \leavevmode\nobreak\ t=1,2,\cdots,k_{1}^{0}(T)\\ \cdots\\ \lambda_{i,m_{0}+1}f_{t}+e_{it}&for\leavevmode\nobreak\ \leavevmode\nobreak\ t=k_{m_{0}}^{0}(T)+1,\cdots,T,\end{cases}$ (1)
where $f_{t}$ is an $r$-dimensional vector of unobserved common factors; $r$ is the number of pseudo-factors in the entire sample; $k_{j}^{0}(T),j=1,\cdots,m_{0}$ are unknown break dates and $\min_{\{j=0,1,\cdots,m_{0}\}}k_{j+1}^{0}(T)-k_{j}^{0}(T)\geq\eta T$ for a constant $0<\eta<1$, with $k_{0}^{0}(T)=0$ and $k_{m_{0}+1}^{0}(T)=T$; $\lambda_{i,j},j=1,\cdots,m_{0}+1$ are factor loadings at time $t$ in regime $j$; and $e_{it}$ is the error term allowed to have serial and cross-sectional dependence as well as heteroskedasticity, and both $N$ and $T$ tend to infinity. $\tau_{j}^{0}=k_{j}^{0}(T)/T\in(0,1),j=1,\cdots,m_{0}$ are break fractions and are fixed constants. For notational simplicity, hereinafter, we suppress the dependence of $k_{j}^{0}$ on $T$. Note that the dimension of $f_{t}$ is the same as that of the pseudo-factors (to be defined soon).
@@
$x_{t}=\begin{cases}\Lambda_{1}f_{t}+e_{t}&for\leavevmode\nobreak\ \leavevmode\nobreak\ t=1,2,\cdots,k_{1}^{0}\\ \cdots\\ \Lambda_{m_{0}+1}f_{t}+e_{t}&for\leavevmode\nobreak\ \leavevmode\nobreak\ t=k_{m_{0}}^{0}+1,\cdots,T,\end{cases}$ (2)
@@
$X_{j}=[x_{k_{j-1}^{0}+1},\cdots,x_{k_{j}^{0}}]^{{}^{\prime}},\leavevmode\nobreak\ \leavevmode\nobreak\ F_{j}=[f_{k_{j-1}^{0}+1},\cdots,f_{k_{j}^{0}}]^{{}^{\prime}},\leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{\boldmath$e$}_{j}=[e_{k_{j-1}^{0}+1},\cdots,e_{k_{j}^{0}}]^{{}^{\prime}}.$
@@
$X=\left[\begin{array}[]{cccccccccc}X_{1}\\ \cdots\\ X_{m_{0}+1}\end{array}\right]$ $=$ $\left[\begin{matrix}F_{1}\Lambda_{1}^{{}^{\prime}}\\ \cdots\\ F_{m_{0}+1}\Lambda_{m_{0}+1}^{{}^{\prime}}\end{matrix}\right]+\left[\begin{array}[]{cccccccccc}\mbox{\boldmath$e$}_{1}\\ \cdots\\ \mbox{\boldmath$e$}_{m_{0}+1}\end{array}\right]=\left[\begin{matrix}F_{1}(\Lambda B_{1})^{{}^{\prime}}\\ \cdots\\ F_{m_{0}+1}(\Lambda B_{m_{0}+1})^{{}^{\prime}}\end{matrix}\right]+\left[\begin{array}[]{cccccccccc}\mbox{\boldmath$e$}_{1}\\ \cdots\\ \mbox{\boldmath$e$}_{m_{0}+1}\end{array}\right],$ (12)
$=$ $\underbrace{\left[\begin{matrix}F_{1}B_{1}^{{}^{\prime}}\\ \cdots\\ F_{m_{0}+1}B_{m_{0}+1}^{{}^{\prime}}\end{matrix}\right]}_{G}\Lambda^{{}^{\prime}}+\left[\begin{array}[]{cccccccccc}\mbox{\boldmath$e$}_{1}\\ \cdots\\ \mbox{\boldmath$e$}_{m_{0}+1}\end{array}\right],$ (16)
@@
denote the numbers of effective factors, i.e., the rank of the common component in regime $k_{j-1}^{0}+1,k_{j}^{0}]$. In (17), we define
@@
as the pseudo-factors because the final expression in (17) provides an observationally equivalent representation, maintaining the structure of the loadings matrix $\Lambda$ unchanged. More precisely, when the break is omitted during the estimation, the factors derived from a full-sample PCA correspond to the pseudo-factors $G$ presented in (17). It is well-established that the presence of breaks can expand the factor space; consequently, for $j=1,\dots,m_{0+1}$, we have $r_{j}\leq r$, where $\text{rank}(G)=r$. This representation allows for changes in factor loadings and the number of factors.
We require the above model to have a minimum representation in the sense that
$rank(\Lambda)=rank(G)=rank([B_{1},B_{2},\dots,B_{m_{0}+1}]).$
The first equality is commonly imposed in the literature on factor model without structural changes, where $B_{j}=I_{r}$ for $j=1,...,m_{0}+1$, rendering the second equality trivial. However, in our context, the second equality is also required, as not all $B_{j}$ matrices necessarily possess full rank $r$. This rank corresponds to the number of pseudo factors for the entire sample.

### v3

@@
Abstract
This paper proposes a quasi-maximum likelihood (QML) estimator for break points in high-dimensional factor models, specifically accounting for multiple structural breaks. We begin by establishing a necessary and sufficient condition to categorize two distinct types of breaks in factor loadings: singular changes and rotational changes. The analysis of the nearly singular subsample covariance matrices of the pseudo-factors plays a key role in our approach. It allows us to demonstrate that the QML estimator precisely identifies the true breakpoint with probability tending to one for singular changes. For rotational changes, we demonstrate that the estimator exhibits stochastically bounded estimation errors, implying break fraction consistency. Furthermore, we introduce an information criterion to estimate the number of breaks, proving that it can detect the true number with probability tending to one. Monte Carlo simulations confirm the strong finite sample performance of our proposed methods. Finally, we provide an empirical example to estimate structural breakpoints in the FRED-MD dataset spanning 1959 to 2024.
Keywords: Quasi-Maximum Likelihood, Singularity, Consistency, Information criterion.
@@ 1 Introduction
Factor models have become a cornerstone in the analysis of large-scale economic and financial datasets characterized by both high cross-sectional and time dimensions. The models’ versatility has led to widespread applications across economics and statistics, including the study of policy shocks (Stock and Watson, 2016; Han, 2025), macroeconomic dynamics (Boivin and Giannoni, 2006), asset pricing (Giglio and Xiu, 2021; Giglio et al., 2025), portfolio theory (Ross, 1976; Chamberlain and Rothschild, 1983; Connor and Korajczyk, 1986, 1988; Ding et al., 2021), causal inference (Gobillon and Magnac, 2016; Xu, 2017), and high-dimensional statistical machine learning (Fan et al., 2021). However, despite their wide applicability, real-world data are rarely stable over time. Structural breaks are ubiquitous, arising from policy interventions, financial crises, technological progress, or global events such as the COVID-19 pandemic. Importantly, they often embody the causal impact of interventions and shocks—policies may reshape economic structures, innovations may alter production and financial linkages, and major crises may transform underlying relationships. Identifying and modeling these breaks is therefore a powerful tool for causal inference, policy evaluation, and understanding how economic systems evolve in response to major events.
@@
Previous studies (e.g., Han and Inoue, 2015) have shown that a single structural break in the factor loading matrix admits an observationally equivalent representation with time-invariant loadings and an expanded set of pseudo-factors. This expansion induces singularity in the covariance matrices of the pseudo-factors from either the pre-break or post-break subsample. Duan, Bai and Han (2023, hereafter DBH) exploited this property and demonstrated that such singularity is key to establishing the consistency of the QML estimator in the single-break setting. However, the DBH approach is restricted by its single-break assumption. In comparison, our framework allows for multiple breaks and establishes a necessary and sufficient condition for break taxonomy. This is a critical advancement, as singularity alone no longer suffices to distinguish break types in multi-break settings. Moreover, we develop a novel proof strategy to overcome the technical challenges unique to multi-break setups and propose an information criterion that consistently selects the number of breaks.
@@
Partly due to the simultaneous existence of both singular and rotational breaks, extending the analysis from a single-break to a multi-break setting presents several significant and non-intuitive challenges. First, unlike the single-break case, merely checking the singularity of the pseudo-factor covariance matrix within a given regime is insufficient to characterize the type of break in a multi-break environment. A break can be rotational even if neither the preceding nor the succeeding regime exhibits full column rank, a phenomenon that can occur due to singular breaks at other time points. Thus, a primary challenge lies in rigorously defining rotational and singular changes in factor loadings within a multi-break framework. Second, because multiple breakpoints are jointly estimated, the estimation errors of these individual breakpoint estimators become inherently interdependent, posing substantial technical challenges for asymptotic analysis. Third, given that the QML estimator exhibits distinct convergence rates for singular and rotational breaks, consistently determining the true number of breaks without prior knowledge of their specific types is challenging.
@@
We address these challenges and make the following contributions. First, we establish a necessary and sufficient condition to precisely distinguish between singular and rotational structural breaks in factor loadings within a multi-break framework. Specifically, for each breakpoint, break types are identified by comparing the numbers of effective factors in the pre-break and post-break regimes with the number of pseudo-factors in the regime formed by pooling these two regimes. Second, to tackle the interdependence of estimation errors mentioned previously, we analyze the asymptotic properties of the $j$-th breakpoint by “conceptually” overfitting the remaining true breakpoints. This strategy requires us to compare the eigenvalues of the sample covariance matrix of the pseudo-factors in a regime against those from a subsample of that same regime. Third, leveraging this crucial result, we prove that the QML estimator identifies the true breakpoint with probability tending to one in large samples for singular breaks, and achieves $O_{p}(1)$ estimation errors for rotational breaks. Finally, we analyze how underfitting and overfitting the number of breaks affects the value of our QML objective function under both types of breaks. This analysis enables us to design an effective penalty function for our information criterion, ensuring its consistent selection of the true number of breaks without requiring prior knowledge of break types.
@@
Links to the literature. A growing body of literature investigates tests for structural breaks in high-dimensional factor models (e.g., Chen et al., 2014; Han and Inoue, 2015; Yamamoto and Tanaka, 2015; Bai et al., 2024; Fan et al., 2024a; Fan et al., 2024b; Peng et al., 2025). Our study differs by focusing on the estimation of breakpoints rather than testing for their presence. Moreover, traditional literature on structural break estimation primarily addresses time series with small cross-sectional dimensions (e.g., Bai and Perron, 1998; Bai, 2000; Qu and Perron, 2007; Chan et al., 2014, 2021). In contrast, our factor model framework operates in a setting where both the cross-sectional ($N$) and time ($T$) dimensions are large.
@@
Within this large $N$, large $T$ context, prior work has explored structural breaks in various settings. Qian and Su (2016), for instance, applied group fused Lasso to panel models with structural breaks, and Li et al. (2016) extended this to panel data with interactive fixed effects. While these studies considered large panels, their focus differed from ours, as they concentrated on breakpoints in regression coefficients rather than the high-dimensional factor loading matrix. For high-dimensional time series, Bai (2010) developed a least squares (LS) estimator for breakpoints in the mean and variance, and Li et al. (2024) introduced a two-way MOSUM procedure with asymptotic properties for change-point estimators in trend functions. Our framework distinguishes itself from these by specifically addressing breakpoints in the loading coefficients of unobserved factors.
@@
Regarding breakpoint estimation in factor models, earlier literature established the break fraction consistency, specifically that $\hat{\tau}-\tau_{0}=o_{p}(1)$, where $\hat{\tau}$ is the estimated breakpoint $\hat{k}$ divided by sample size $T$, and $\tau_{0}$ is the true fraction (e.g., Chen, 2015; Cheng et al., 2016). More recent studies have provided theoretical results on the detailed rate of the error bound for the estimated breakpoint. For example, Barigozzi et al. (2018) proposed a wavelet-based method to estimate both the number and locations of structural breaks in common and idiosyncratic components. Barigozzi et al. (2025) further developed a MOSUM procedure for detecting multiple change points in high-dimensional factor models. Both studies showed that the estimation error of the breakpoint diverges slowly with sample size, implying a consistently estimated break fraction. Similarly, Baltagi et al. (2017, 2021) developed an estimator that compares subsample second moments of factors estimated via full-sample principal components. Their breakpoint estimator exhibits a stochastically bounded estimation error, which also implies a consistently estimated break fraction.
@@
In contrast to these approaches, our QML estimator achieves not only fraction-consistency but also point-consistency (sometimes referred to as super-consistency in the literature), meaning the estimated breakpoints equal the true ones with probability approaching one in large samples. To our knowledge, only a few studies have established point-consistency for breakpoint estimators of the loading matrix in factor models. Bai et al. (2020) demonstrated that the LS estimator is point-consistent under both large and small break settings. However, their framework is limited to a single break, whereas our methodology accommodates multiple breaks. Ma and Su (2018) proposed a fused Lasso estimator that is consistent for multiple breakpoints in factor models, under the assumption of a constant number of factors. Our work considers a more general and empirically relevant setting by allowing the number of effective factors to vary over time. Additionally, Ma and Tu (2023) adopted a group fused Lasso approach to estimate changes in regression coefficients of one factor on the remaining factors. Their fused Lasso estimator exhibits stochastically bounded errors, which are then refined by LS to achieve point-consistency. Nonetheless, their method does not accommodate certain types of rotational changes (e.g., when the post-break loading matrix is rescaled by a constant), a feature that our Type 2 break in Section 3 does allow.
@@
Notation. The following notation is used throughout the paper. Let $\rho_{i}(\mathbb{B})$ denote the $i$-th eigenvalue of an $n\times n$ symmetric matrix $\mathbb{B}$, and $\rho_{1}(\mathbb{B})\geq\rho_{2}(\mathbb{B})\geq\cdots\geq\rho_{n}(\mathbb{B})$. For an $m\times n$ real matrix $\mathbb{A}$, we denote its Frobenius norm as $\|\mathbb{A}\|=[tr(\mathbb{A}\mathbb{A}^{{}^{\prime}})]^{1/2}$, and its adjoint matrix as $\mathbb{A}^{\#}$ when $m=n$. Let $\mathrm{Proj}(\mathbb{A}|\mathbb{Z})$ denote the projection of columns of matrix $\mathbb{A}$ onto the columns of matrix $\mathbb{Z}$. For a,b>0a,\ b>0, $a\gtrsim b$ indicates that there exists a constant $c>0$ such that $a\geq cb$. We write $a\asymp b$ to mean that there exist constants $c_{1},c_{2}>0$ such that $c_{1}b\leq a\leq c_{2}b$. For a real number $x$, $[x]$ represents the integer part of $x$.
@@ 2.1 Model
We consider the following factor model with $m_{0}$ common breaks at $k_{1}^{0}(T)<k_{2}^{0}(T)<\cdots<k_{m_{0}}^{0}(T)$ in the factor loadings for $i=1,\cdots,N$:
$x_{it}=\begin{cases}\lambda_{i,1}^{\prime}f_{t}+e_{it}&for~~t=1,2,\cdots,k_{1}^{0}(T)\\ \cdots\\ \lambda_{i,m_{0}+1}^{\prime}f_{t}+e_{it}&for~~t=k_{m_{0}}^{0}(T)+1,\cdots,T,\end{cases}$ (1)
where $f_{t}$ is an $r$-dimensional vector of unobserved common factors; $r$ is the number of pseudo-factors in the entire sample; $k_{j}^{0}(T),j=1,\cdots,m_{0}$ are unknown break dates and $\min_{\{j=0,1,\cdots,m_{0}\}}k_{j+1}^{0}(T)-k_{j}^{0}(T)\geq\eta T$ for a constant $0<\eta<1$, with $k_{0}^{0}(T)=0$ and $k_{m_{0}+1}^{0}(T)=T$; $\lambda_{i,j}$ are $r\times 1$ factor loading vectors for unit $i$ in the $j$th regime $t\in[k_{j-1}^{0}(T)+1,\,k_{j}^{0}(T)]$ for $j=1,\cdots,m_{0}+1$; and $e_{it}$ is the error term allowed to have serial and cross-sectional dependence as well as heteroskedasticity, and both $N$ and $T$ tend to infinity. $\tau_{j}^{0}=k_{j}^{0}(T)/T\in(0,1),j=1,\cdots,m_{0}$ are break fractions and are fixed constants. For notational simplicity, hereinafter, we suppress the dependence of $k_{j}^{0}$ on $T$. Note that the dimension of $f_{t}$ is the same as that of the pseudo-factors (defined below).
@@
$x_{t}=\begin{cases}\Lambda_{1}f_{t}+e_{t}&for~~t=1,2,\cdots,k_{1}^{0}\\ \cdots\\ \Lambda_{m_{0}+1}f_{t}+e_{t}&for~~t=k_{m_{0}}^{0}+1,\cdots,T,\end{cases}$ (2)
@@
$X_{j}=[x_{k_{j-1}^{0}+1},\cdots,x_{k_{j}^{0}}]^{{}^{\prime}},~~F_{j}=[f_{k_{j-1}^{0}+1},\cdots,f_{k_{j}^{0}}]^{{}^{\prime}},~~\mbox{$e$}_{j}=[e_{k_{j-1}^{0}+1},\cdots,e_{k_{j}^{0}}]^{{}^{\prime}}.$
@@
$X=\left[\begin{array}[]{cccccccccc}X_{1}\\ \cdots\\ X_{m_{0}+1}\end{array}\right]$ $=$ $\left[\begin{matrix}F_{1}\Lambda_{1}^{{}^{\prime}}\\ \cdots\\ F_{m_{0}+1}\Lambda_{m_{0}+1}^{{}^{\prime}}\end{matrix}\right]+\left[\begin{array}[]{cccccccccc}\mbox{$e$}_{1}\\ \cdots\\ \mbox{$e$}_{m_{0}+1}\end{array}\right]=\left[\begin{matrix}F_{1}(\Lambda B_{1})^{{}^{\prime}}\\ \cdots\\ F_{m_{0}+1}(\Lambda B_{m_{0}+1})^{{}^{\prime}}\end{matrix}\right]+\left[\begin{array}[]{cccccccccc}\mbox{$e$}_{1}\\ \cdots\\ \mbox{$e$}_{m_{0}+1}\end{array}\right],$ (12)
$=$ $\underbrace{\left[\begin{matrix}F_{1}B_{1}^{{}^{\prime}}\\ \cdots\\ F_{m_{0}+1}B_{m_{0}+1}^{{}^{\prime}}\end{matrix}\right]}_{G}\Lambda^{{}^{\prime}}+\left[\begin{array}[]{cccccccccc}\mbox{$e$}_{1}\\ \cdots\\ \mbox{$e$}_{m_{0}+1}\end{array}\right],$ (16)
@@
denote the numbers of effective factors in regime $j$, i.e., the rank of the common component in regime $k_{j-1}^{0}+1,k_{j}^{0}]$. In (17), we define
@@
as the pseudo-factors because the final expression in (17) provides an observationally equivalent representation, maintaining the structure of the loadings matrix $\Lambda$ unchanged. More precisely, when the break is omitted during the estimation, the factors derived from a full-sample PCA correspond to the pseudo-factors $G$ presented in (17). It is well established that the presence of breaks can expand the factor space; consequently, for $j=1,\dots,m_{0}+1$, we have $r_{j}\leq r$, where $\text{rank}(G)=\text{rank}([B_{1},B_{2},\dots,B_{m_{0}+1}])=r$. This representation allows for changes in factor loadings and the number of factors. Concrete constructions of $\Lambda$, $B_{j}$, $r$, and $r_{j}$ under different breakpoint types are given in Examples 1–3 of Section 3. We differentiate between effective factors and pseudo-factors. Effective factors are the factors required to capture common variation within a single, stable regime. In contrast, pseudo-factors are used to describe the factors (whose dimension may exceed that of effective factors due to breaks) within either the full sample (i.e., $G$) or a subsample consisting of consecutive regimes.